OpenStudy (anonymous):

A storage tank contains 21.6kg of nitrogen (N2)at an absolute pressure of 3.65 atm. what will be the pressure if nitrogen is replaced by an equal mass of CO2 or carbon dioxide?..thanks!

OpenStudy (anonymous):

A little less. The pressure of a gas is determined by the number of particles in it (recall the ideal gas equation of state, p = nRT/V, where n = number of moles of the gas). Since each CO2 molecules is heavier (44.0 Da) than each N2 molecule (28.0 Da), replacing 100 g of N2 with 100 g of CO2 means replacing a certain number of N2 molecules with FEWER CO2 molecules. Fewer molecules => lower pressure.

OpenStudy (anonymous):

i understand that much. im not sure how to do the equation

OpenStudy (anonymous):

I see. OK, we're not going to be able to get a numerical answer, because you haven't said how much N2 is to be replaced. But we can get an equation for the anwer, sure.

OpenStudy (anonymous):

im assuming its the whole 21.6 kg of nitrogen that is being replaced

OpenStudy (anonymous):

Oh, that's easy then.

OpenStudy (anonymous):

\[\frac{p_2}{p_1} = \frac{n_2}{n_1} = \frac{m / A_C}{m / A_N} = \frac{A_N}{A_C} = \frac{28.0}{44.0}\]

OpenStudy (anonymous):

awesome. so that is the ratio of the molecular masses, correct?

OpenStudy (anonymous):

how do i know wat the volume is? is that just the STP of 22.4L?

OpenStudy (anonymous):

You don't need to know the volume.

OpenStudy (anonymous):

then how do i find out the pressure if im using the ideal gas law ?

OpenStudy (anonymous):

we got the moles bu i still dont know how to get a numerical value forthe pressure

OpenStudy (anonymous):

Look again at my equation. The ratio of moles equals the ratio of pressures. You know the initial pressure, p1.

OpenStudy (anonymous):

i still dont understand how that answers the initial question =/ im sorry

OpenStudy (anonymous):

p1 = pressure if all N2 = 3.65 atm p2 = pressure of all CO2 p2 = (p2/p1) * p1 p2/p1 = A_N2/A_CO2 = 28.0/44.0